OpenStudy (anonymous):

Identify the first step using the elimination method to solve the following system of equations: 3x + 2y = 4 -5x + 15y = -25 A. 5(3x + 2y = 4) 3(-5x + 15y = -25) B. 15(3x + 2y = 4) 2(-5x + 15y = -25) C. 3(3x + 2y = 4) 5(-5x + 15y = -25) D. 25(3x + 2y = 4) 4(-5x + 15y = -25)

hero (hero):

@AntoinetteWillman , which do you believe is correct?

OpenStudy (anonymous):

B.

OpenStudy (anonymous):

@Hero

hero (hero):

No, it's not B unfortunately.

hero (hero):

You have to multiply each equation by a number such that one of the variables can be eliminated by addition or subtraction.

hero (hero):

For example, if you had the following system: -2x + 3y = 4 3x + 4y = 5 Then you could multiply the first equation by 3 3(-2x + 3y = 4) And the second equation by 2: 2(3x + 4y = 5) As a result, you will get: -6x + 9y = 12 6x + 8y = 10 Now notice that you have -6x in the first equation and 6x in the second equation. Because -6x + 6x = 0, the x variable is eliminated when adding both equations together.

hero (hero):

@AntoinetteWillman

OpenStudy (anonymous):

So C.

hero (hero):

No, not C either. I suppose you're going to keep guessing.

OpenStudy (anonymous):

I give up. Thanks anyway.

hero (hero):

I'm sorry to hear that.